OpenStudy (anonymous):

Determine if triangle RST with coordinates R (3, 4), S (5, 5), and T (6, 1) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific.

OpenStudy (anonymous):

@Hero @nincompoop @robtobey @Compassionate

hero (hero):

Hint: https://www.desmos.com/calculator/fq8bdv3qo6

OpenStudy (anonymous):

No, it's not a right triangle, it's an Isosceles. @Hero

hero (hero):

Isosceles? Who said it was isosceles? https://www.desmos.com/calculator/kp5ypkg6y6

OpenStudy (anonymous):

It looks like an isosceles triangle. Instead of a right triangle.

hero (hero):

In mathematics, you verify things using mathematical law and reasoning, because looks can be deceiving.

OpenStudy (anonymous):

I'm still not sure how to solve the problem, though..@Hero

OpenStudy (anonymous):

@phi

OpenStudy (phi):

there a two ways to prove you have (or do not have) a right triangle: 1) show it has a right angle, or 2) show pythagoras works

OpenStudy (anonymous):

I'm still not sure. @phi

OpenStudy (anonymous):

It doesn't have a right angle, from what I can see.

OpenStudy (phi):

to show it has a right angle, we use the idea that perpendicular lines have slopes that are *negative reciprocals* for example, 2 and -½ , or ¾ and -4/3 (flip and then multiply by -1 to go from one slope to the other) that means we need to find the slopes between points R and S S and T T and R and see if any pair is the negative reciprocal of the other

OpenStudy (phi):

can you find the slope between R and S R(3, 4), S (5, 5) ?

OpenStudy (anonymous):

\[\frac{ 3-5 }{ 4-5 } = \frac{ -2 }{ -1}\]

OpenStudy (phi):

close. but it is change in y divided by change in x you did change in x divided by change in y

OpenStudy (anonymous):

So, it'd be \[\frac{ -1 }{-2 }\]

OpenStudy (phi):

yes. and minus divided by minus is plus so just 1/2

OpenStudy (anonymous):

okay. Then what?

OpenStudy (phi):

now let's find the slope between S and T

OpenStudy (anonymous):

S = (5, 5) T = (6, 1) \[\frac{ 5 - 1 }{ 5 - 6 } = \frac{ 4}{ -1}\]

OpenStudy (phi):

yes. can you simplify that ?

OpenStudy (anonymous):

I don't know how..

OpenStudy (phi):

it has only 1 minus sign so the answer is negative in other words \[ \frac{4}{-1}= - \frac{4}{1} \] now what is 4 divided by 1 ?

OpenStudy (anonymous):

4.

OpenStudy (phi):

so the answer is -4 so far we have slope R to S = ½ slope S to T = -4 those are not negative reciprocals so the angle formed by RS and ST is not a right angle. one more side to try find the slope between R and T

OpenStudy (anonymous):

R = (3, 4) T = (6, 1) \[\frac{ 4-1 }{ 3-6 } = \frac{ 3 }{ -3}\]

OpenStudy (phi):

yes. now simplify it

OpenStudy (anonymous):

Is it 1 ?

OpenStudy (phi):

it has a minus sign (in the bottom) so the answer will be negative. you get -1

OpenStudy (phi):

here is what you know negative reciprocal slope R to S = ½ -2 slope S to T = -4 1/4 slope R to T = -1 1 there is no pair that form negative reciprocals, so you know none of the lines are perpendicular to each other. In other words, none of the lines cross at 90º so this triangle does not have a 90 degree angle, so it is *not* a right triangle

OpenStudy (phi):

Now for the hard part. If it is not a right triangle, what changes can be made to make it a right triangle?

OpenStudy (phi):

Our best bet is notice that slope R to T = -1 the negative reciprocal is +1 so I would move point S so that the slope from R to S is +1 that will make RS and RT perpendicular

OpenStudy (phi):

so far we have R(3, 4), S (5, 5) the slope is \[ \frac{5-4}{5-3} \] any idea how to change S so that we get +1 ?

OpenStudy (phi):

to get a 1, you want the top number and the bottom number to be the same, because anything divided by itself is 1 if we leave the top number 5-4=1 then you want the bottom number to be 1 we want to change 5-3 to n-3 where n is a new number, so that we get 1 any idea what n should be ?

OpenStudy (anonymous):

@phi i need your help explaining the rest of this i got everything else so far!:)

OpenStudy (anonymous):

@iambatman could you please help me on this?

OpenStudy (anonymous):

@e.mccormick could you please help me on this?